OpenStudy (anonymous):

Select True or False. The expression 3x − 4(2x − 5) is equivalent to the expression 20 − 5x. @Princer_Jones

OpenStudy (anonymous):

3x − 4(2x − 5)=3x-8x+20=20-5x TRUE

OpenStudy (anonymous):

@geny55

OpenStudy (anonymous):

thx :)